OpenStudy (anonymous):

Find the equation of the plane in the xyz-space through the point P=(3,4,2) and perpendicular to the vector n = (-5,-5,5) z = ________

OpenStudy (amistre64):

yay were back lol

OpenStudy (anonymous):

The equation of the plane is the dot product of P and n.

OpenStudy (amistre64):

they give you the normal; and a point, so all we do is attach it all

OpenStudy (anonymous):

ya...i'm aware...i'm getting marked wrong though

OpenStudy (anonymous):

Could you post some of your workings?

OpenStudy (anonymous):

i get it as -5(x-3) -5(y-4) + 5(z-2)

OpenStudy (amistre64):

-5<1,1,-1> (x-Px)+(y-Py)-(z-Pz)=0

OpenStudy (amistre64):

your right, but reduce it

OpenStudy (anonymous):

right...distribute i get -5x -5y + 5z = -25

OpenStudy (amistre64):

and they might not like a -leading coeff

OpenStudy (anonymous):

Looks good to me. You can make it simpler though.

OpenStudy (anonymous):

well i keep getting it marked wrong and i cant figure out why

OpenStudy (amistre64):

format, thats all

OpenStudy (anonymous):

it wants the z = obv.....so if i plug in 2 arbitrary constants for x and y

OpenStudy (amistre64):

x+y-z-5=0 is what i would put for it

OpenStudy (anonymous):

hang on lemme try that

OpenStudy (amistre64):

err.... z = x+y-5 :)

OpenStudy (anonymous):

I would say to rearrange it to be z = x + y -5.

OpenStudy (anonymous):

yes that worked

OpenStudy (anonymous):

u guys are awesome thanks a ton

OpenStudy (amistre64):

yep :)